OpenStudy (janu16):

Given the point (4,5) and the slope of 6, find y when x=24.

OpenStudy (janu16):

do you know what formula to use @DanJS

OpenStudy (danjs):

yes, do you recall y = mx+ b slope = m

OpenStudy (janu16):

ya

OpenStudy (danjs):

they give you a point (4 , 5), x=4 and y=4 put those in also

OpenStudy (danjs):

y=5 *

OpenStudy (danjs):

what is b, the y-intercept value

OpenStudy (janu16):

5

OpenStudy (danjs):

y = mx+b 5 = 6*4 + b

OpenStudy (janu16):

so isnt y and b the same thing?

OpenStudy (danjs):



OpenStudy (danjs):

the y-axis intercept

OpenStudy (danjs):

so the line is y = m*x + b you given m=6 , x = 4, y = 5 only thing left is find b

OpenStudy (danjs):

5 = 6*4 + b 5 = 24 + b b= 5 - 24 b = -19

OpenStudy (danjs):

so now with the slope m=6 and y-int b = -19 you can write the general equation for the line... y = mx+b y = 6x - 19

OpenStudy (janu16):

so so that would be the equation y = 6x - 19

OpenStudy (danjs):

yes, now you can answer the question

OpenStudy (janu16):

118 144 125 149 these are the answer choices so i plug them n or soething?

OpenStudy (danjs):

put in 24 for x

OpenStudy (janu16):

y=6+24-19

OpenStudy (janu16):

nvbm

OpenStudy (janu16):

so it would be 6 times 24 than -19?

OpenStudy (danjs):

yes

OpenStudy (janu16):

omg thanks i finally got it. so like there was a question where they gave you a point and you have to tell which graph the point belongs to. i have never learned tht and its weird

OpenStudy (danjs):

ok open a new thread , this is long

OpenStudy (janu16):

ok

OpenStudy (danjs):

i got time for one more at the moment